OpenStudy (anonymous):

which function has an amplitude of 3 and a period of pi/2

OpenStudy (jack1):

equation of a function \(\large f(x) = a \sin(bx−c)+d\) \( \large a = amplitude\) \( \Large \frac {2\pi}{b} = period\) \( \Large \frac {c}{b} = phase~shift\) \( \large d = vertical~shift\) can you solve from here @vintagelover143 ?